A valószínűségszámítás II. gyakorlaton felajánlott jegyek diákok:

Ágfalvi Veronika  3  közepes
Belicza Zsuzsanna  3  közepes
Bíró Anna  1  pótdolgozat
Bucsay Balázs  2  elégséges

Kereskényiné,

Balogh Zsófia  4  jó
Mikó Tamás  1  pótdolgozat
Petrányi Gábor  2  elégséges
Rauf Bettina  5  jeles
Rábai Zsolt  5  jeles
Varga Nóra  4  jó